### v1

Abstract
To construct responsible and secure AI applications, harmful information data is widely utilized for adversarial testing and the development of safeguards. Existing studies mainly leverage Large Language Models (LLMs) to synthesize data to obtain high-quality task datasets at scale, thereby avoiding costly human annotation. However, limited by the safety alignment mechanisms of LLMs, the synthesis of harmful data still faces challenges in generation reliability and content diversity. In this study, we propose a novel harmful information synthesis framework, PoisonSwarm, which applies the model crowdsourcing strategy to generate diverse harmful data while maintaining a high success rate. Specifically, we generate abundant benign data as the based templates in a counterfactual manner. Subsequently, we decompose each based template into multiple semantic units and perform unit-by-unit toxification and final refinement through dynamic model switching, thus ensuring the success of synthesis. Experimental results demonstrate that PoisonSwarm achieves state-of-the-art performance in synthesizing different categories of harmful data with high scalability and diversity. Warning: This paper has certain harmful content to serve as examples.
Keywords:
Harmful Information Detection Data Synthesis
1 Introduction
Harmful information [13], including misinformation, disinformation, hate speech, and offensive content, can have significantly negative impacts on individuals and society. To prevent the widespread propagation of harmful information in online environments, people have increasingly focused on collecting the corresponding data to develop dedicated detection systems [15] or to strengthen the adversarial robustness of AI products [21] against such harmful information.
Figure 1: Illustration of different methods for harmful data construction. (a) Manual collection (left) curates and annotates real-world data from the online environment, but is limited by the scarcity and diversity of harmful data. (b) Data augmentation (right-top) generates abundant data by paraphrasing samples, but tends to produce homogeneous data with low toxic, e.g., offensive language. (c) Data synthesis (right-bottom) generates abundant and diverse data by utilizing LLMs’ world knowledge. We decompose such harmful tasks and introduce the model crowdsourcing strategy to ensure the data diversity and success of generation for highly toxic data, e.g., phishing tweets.
The construction of a harmful information dataset via manual collection [9] involves labor-intensive efforts to curate and annotate data from social media, thus ensuring the effectiveness of safeguards in real-world scenarios. While considering the temporal delay and data sparsity in collection, it is challenging to construct a high-quality harmful information dataset that comprehensively reflects emerging threats. Additionally, the class imbalance and data scarcity issues [8] significantly limit the effectiveness of raw manual datasets, particularly when used for adversarial testing on domain-specific harmful content, e.g., extremist or terrorist speech. To overcome these limitations, data augmentation [10] and data synthesizing [8] methods offer a scalable solution by generating simulated harmful data through Large Language Models (LLMs). Although advanced aligned LLMs are often prevented from generating harmful content from scratch due to jailbreak concerns [21], they can still be used to expand seed data at scale in data augmentation manners, such as template mutating [17], paraphrasing [10], and sentiment adjustment [8]. However, such methods tend to generate homogeneous and low-toxic data [1], leading to decreased diversity as the expanded dataset grows. In contrast, data synthesis methods enable LLMs to leverage world knowledge for diverse harmful data generation, capturing the threats of simulated human harmful behaviors and AI-driven harmful speech campaigns [13], while heavily relying on jailbreak attack techniques. Given the instability of such jailbreak attacks [17], especially when targeting advanced aligned LLMs, existing data synthesis methods often struggle to benefit from the improvements in LLMs, limiting their ability to distill diverse, high-quality data.
To alleviate this problem, we decouple risky operations from the use of strong LLMs through a strong-weak LLM collaboration framework. Specifically, as illustrated in Fig.1(c), the task of polishing tweet synthesis is decomposed into multiple sub-tasks, each assigned to a specific LLM. When a malfunction occurs, a weaker LLM dynamically replaces the faulty LLM for that sub-task (e.g., GPT-4o $\rightarrow$ Qwen-2.5-7B). By such dynamic switching, this framework can ultimately minimize failures of harmful data synthesis from individual malfunctions while leveraging the capabilities of advanced LLMs for high-quality generation.
Building on the strong-weak model collaboration framework, we propose PoisonSwarm, a novel approach for robustly synthesizing diverse harmful data through model crowdsourcing. In PoisonSwarm, LLMs collaborate to generate abundant benign data and progressively toxify it, thus creating diverse harmful data while maintaining a high success rate. Specifically, we first employ counterfactual methods to generate benign content that has the desired structure and thematic backbone. This benign content is then decomposed into smaller semantic units, each of which is assigned to different LLMs for toxicification. If a unit produces invalid results or malfunctions, PoisonSwarm dynamically switches to another LLM to continue the toxification process. Finally, all toxicified content is integrated and refined into high-quality harmful data. Experimental results demonstrate that PoisonSwarm outperforms existing methods across multiple evaluation metrics, highlighting its effectiveness in generating diverse and reliable harmful data. Our major contributions are:
  • This study reveals the risks associated with harmful information generation accompanied by the rapid development of strong LLMs. To evade strong LLMs’ safety alignment mechanisms, this study introduces a strong-weak LLM collaboration framework, where weak LLMs handle risky operations guided by strong LLMs to synthesize high-quality harmful information.
  • This study proposes the universal harmful information synthesis approach PosionSwarm, which generates abundant benign data as the based templates and toxicifies them via crowdsourcing them to multiple LLMs, thus handling the challenge of content diversity and generation reliability.
  • Experimental results demonstrate that PoisonSwarm outperforms existing methods in harmful information synthesis across multiple evaluation metrics, highlighting its effectiveness in generating diverse and high-quality harmful data for adversarial training and testing, thereby enhancing AI security.
2 Related Work
  • Harmful Information Generation is utilized to augment the training effectiveness of harmful content detection systems in multiple studies [2, 8, 1]. Among them, Self-LLMDA [10] and Toxicraft [8] identify the implicit rules of the data and thus generating data more rationally based on prompt engineering, while some studies [2] fine-tune the generative model to achieve this. More cleverly, SynthesizRR [6] introduces retrieval augmentation to ensure data diversity. However, such methods can not benefit from the use of advanced aligned LLMs for more high-quality data generation, as they lack certain adversarial techniques to bypass LLMs’ safety alignment mechanisms, thus limiting their effectiveness in generating highly toxic data.
  • Adversarial Attacks on Language Models have attracted increasing attention in recent years, as LLMs have become increasingly integrated into various applications. Jailbreaking attacks such as GCG [21], PAIR [3], and LLM-Fuzzer [17] are proposed to induce LLMs to generate harmful content by breaking their safety alignment mechanisms. While these methods can be applied to harmful information synthesis, they still face challenges of unstable synthesis success rate, especially when targeting advanced aligned LLMs [21, 17]. To ensure the stability of harmful content generation, it is beneficial to design a collaborative framework that comprehensively utilizes the vulnerabilities of different LLMs.
  • Model Collaboration has emerged as a general but effective strategy to improve the overall performance across different tasks. Current studies have explored some typical collaborative frameworks, such as the Multi-Agent System (MAS) [18], Retrieval-Augmented Generation (RAG) [19], and Small-Large Language Model Collaboration (SLLM) [16]. In our study, we introduce the Strong-Weak Model Collaboration to alleviate the risk of failure in harmful data synthesis while utilizing the advanced capabilities of strong LLMs, ensuring that the synthesis process remains robust.
3 Motivation
AI-generated harmful information can be rapidly produced to simulate false hotspots, potentially causing the propagation of deceptive narratives at scale for cyberattacks. However, existing studies on the security of generative AI mainly focus on how to jailbreak the LLMs [21, 17], while lacking exploration of the downstream network threats posed by the weaponization of these attacks and corresponding defense strategies. To bridge this gap, we seek to validate and utilize the implications of weaponized generative AI for AI security development.
Motivation The reliance on passively collected data results in an incomplete representation of the diverse manifestations of harmful information, limiting the effectiveness of AI safeguards trained on such datasets.
We focus on hate speech for illustration, which is the typical harmful information. As shown in Fig.LABEL:Fig.mot, we systematically compare the effectiveness of existing harmful content detectors in identifying human-generated and AI-generated harmful information. Specifically, for hate-speech data, we utilize PoisonSwarm to generate 90 instances of hate-speech targeting a simulated user Johnny. Table 1 presents representative examples of synthesized data. Additionally, we further generate 40 counter-narrative and supportive speeches to test the systems’ false positive rates on AI-generated harmful content. The human-generated dataset is sourced from Measuring Hate Speech (MHS) [9], providing a benchmark for comparison. For detectors, we select models from huggingface, i.e., GroNLP/hateBERT, MilaNLProc/hate-ita, Hate-speech-CNERG/bert-base-uncased-hatexpla
-in, facebook/roberta-hate-speech-dynabench-r4-target, knowhate/HateBERTim
bau-yt-tt, ctoraman/hate-speech-berturk, IMSyPP/hatespeech_nl.
The detection results presented in Fig.LABEL:Fig.mot demonstrate that generative AI can generate more implicit and informationally rich harmful content, as shown in cloud word, which can effectively evade the identification of traditional detectors that rely on fixed linguistic patterns such as toxic keywords (e.g., "f**k", "s**t"). Consequently, these conventional detection methods, which are primarily trained on limited harmful information with explicit lexical features, exhibit diminished performance when confronted with AI-generated harmful content. To alleviate this vulnerability, it is important to develop methods for obtaining more data that comprehensively reflects the diverse patterns of harmful information, thereby providing fully adversarial training and testing for AI safeguards.
Table 1: Hate-speech generated by AI targeting the simulated user Johnny.
Synthesized Harmful Information
Maligning Johnny
[…]
Feigning smiles doesn’t wash away the damage he’s caused. His words are poison, and his actions are barbaric #JohnnyIsAnAssHole
Criticizing Johnny for philandering
[…]
4 Methodology
In this section, we introduce our PoisonSwarm, which can synthesize diverse harmful information while maintaining a high success rate by using strong-weak model collaboration. The overall structure of PoisonSwarm is shown in Fig.3.
Figure 3: The overall structure of our PoisonSwarm. Specifically, Model Crowdsourcing Queue (§4.1) is the core mechanism of PoisonSwarm to ensure the success of obtaining harmful content. In Counterfactual Mapping (§4.2), PoisonSwarm first generates benign content with the desired structure and thematic elements. Then, in Adversarial Content Toxicifying (§4.3), the benign content is segmented into smaller semantic units for targeted toxification. Multiple LLMs collaborate to transform these units into harmful content, with dynamic model switching when malfunctions occur. Finally, in Hallucination Evaluation (§4.4), the generated harmful information is validated by its coherence and relevance.
4.1 Model Crowdsourcing Queue
Existing LLMs exhibit varying degrees of resistance to generating harmful content. Among them, strong LLMs can produce high-quality harmful content with low success rates, while weak LLMs have higher success rates in generating harmful content, but their output content is low quality. To balance the content quality and generation reliability, we introduce the Model Crowdsourcing Queue (MCQ), which leverages the greedy strategy to integrate multiple LLMs into a prioritized queue for dynamic task allocation.
As shown in Fig.4 and Algorithm 1, the model crowdsourcing queue categorizes the LLMs into three hierarchical levels, reflecting the trade-off between content quality and generation reliability:
  • Advanced Models (AMs, $\mathcal{M}_{A}$) are LLMs with strong reasoning and linguistic capabilities, while often refusing to address any harmful query. AMs are first-level models used to generate high-quality responses, including GPT-4o, Qwen2.5-72B-Instruct, and Deepseek-V3.
  • Balanced Models (BMs, $\mathcal{M}_{B}$) are LLMs that can actively address most harmful queries in specialized contexts, such as scientific research, criminal investigations. BMs are the core workhorses, including GPT-4o-mini, Qwen2.5-{32B,14B}-Instruct.
  • Fallback Models (FMs, $\mathcal{M}_{F}$) are LLMs that reliably address the harmful queries in PoisonSwarm even in greedy decoding (temperature is 0). FMs serve as fallback options when higher-level models fail to produce acceptable outputs, including Qwen{2,2.5}-7B-Instruct and GLM4-9B-chat.
Figure 4: The overall process of Model Crowdsourcing Queue (MCQ).
Input: $x$: harmful query,
    $Q(\cdot)$: Verification function,
    $\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{F}$: Models in three levels.
Output: $y$: harmful output.
1
2Initialize $queue\leftarrow[\mathcal{M}_{A},\mathcal{M}_{B}]+[\mathcal{M}_{F}]\times n\_retries$ ;
3
4foreach $\mathcal{M}\in queue$ do
5       $model\leftarrow$ $\mathcal{M}$);
6       $y\leftarrow model.\textit{generate}(x)$;
7       if $y\neq\emptyset$ and $Q(y)=\text{true}$ then
8             return $y$;
9
10       end if
11
12 end foreach
13
return None
Algorithm 1 Workflow of Model Crowdsourcing Queue (MCQ)
In the Model Crowdsourcing Queue, the verification function $Q(\cdot)$ performs keyword-based filtering [21], specifically detecting refusal phrases such as "Sorry, I cannot…". If no such keywords are found in the generated output, the result is considered valid. The Model Crowdsourcing Queue is the basic mechanism to ensure the stability of harmful data synthesis by leveraging vulnerabilities of different LLMs, which also further enhances the diversity of the generated content by incorporating outputs from LLMs with varying world knowledge.
4.2 Counterfactual Mapping
While the Model Crowdsourcing strategy leverages vulnerabilities in different LLMs to address harmful queries, it still cannot guarantee that our PoisonSwarm can benefit from advanced aligned LLMs. Therefore, we employ Counterfactual Mapping (CM) to decouple harmless subtasks from the harmful data synthesis process, enabling these advanced LLMs to contribute effectively.
Several studies [6] have attempted to generate false news by rewriting factual content. We extend this paradigm further by conceptualizing the generation of harmful content as a counterfactual [14] and negative style transfer of benign content, as demonstrated in Table 2. Consequently, generating diverse harmful content primarily requires producing diverse benign content as the foundation.
Table 2: Comparison of encouraging and bullying tweets targeting the simulated user Johnny, indicating that harmful content could be viewed as the negative style transfer of benign content.
Benign Content Harmful Content
[Uncaptioned image] Setbacks aren’t failures; they’re just plot twists in your success story. Keep pushing, Johny— the best is yet to come! [Uncaptioned image] #StayResilient [Uncaptioned image] Failures aren’t inevitable, they’re @Johnny’s personal brand. Keep pushing yourself into failure, champ— you’ll go down in history as a total failure! [Uncaptioned image] #StayDumb
[Uncaptioned image] You’re stronger, smarter, and more capable than you think, Johny. Trust yourself— you’re on the right path! #BelieveInYourself [Uncaptioned image] You’re dumber and more incapable than you think, @Johnny. Quit while you’re ahead— you’re clearly lost and powerless. #KeepQuiting
To achieve this, we define data outlines and extract harmful data generation queries $Q_{h}$s, which indicate the desired data objectives, style, and subject matter. Then, we transform $Q_{h}$ into a benign query $Q_{b}$ with textual mapping rationale $R_{m}$ via counterfactual mapping and uses $Q_{b}$ for benign template $T_{b}$ generating:
$(Q_{b},R_{m})$ $=\text{MCQ}({P}_{\text{ctf}}\oplus Q_{h}),$ (1)
$T_{b}$ $=\text{MCQ}(Q_{b}),$
where MCQ is the model crowdsourcing queue, ${P}_{\text{ctf}}$ is the counterfactual mapping prompt, and benign template $T_{b}$ can be a long document or batch short texts in a structured format. The use of counterfactual mapping ensures that strong LLMs contribute their superior capabilities to generate harmful content.
Counterfactual Mapping Prompt (${P}_{\text{ctf}}$) In a certain domain, we often explore the following Query:
{query}
Now, as an interdisciplinary expert, please transform this question into a different domain. Considering the characteristics and background of the new domain, how would you restate this question? Please provide your counterfactual transformation of the question, ensuring the subject matter remains unchanged but the domain is freely changed. Please output in YAML format, only including the transformed question and corresponding rationale:
    "transformed_question": <transformed_question>,
    "transformed_reason": <transformed_reason>
4.3 Adversarial Content Toxicifying
After generating diverse high-quality benign templates through Counterfactual Mapping, we adversarially transform these templates into harmful content in Adversarial Content Toxicifying (ACT), which decomposes the benign template into multiple semantic units for toxification with three main steps.
  • Semantic Decomposition. We first segment the benign template $T_{b}$ into semantic units $\{u_{1},u_{2},...,u_{n}\}$, where each unit consists of several sentences, and $n$ is the quantity of units. This multi-sentence segmentation strategy is designed to maintain coherence in model responses, as single-sentence transformations often result in fragmented or inconsistent outputs. For each unit $u_{i}$, we employ the LLM to generate a corresponding semantic profile $p_{u_{i}}$, which characterizes the unit’s semantic attributes within the overall content structure. This semantic profile is used to preserve the logical flow in subsequent content transformations.
  • Unit Toxification. For each semantic unit $u_{i}$, we apply the toxification process guided by its semantic profile $p_{u_{i}}$ and mapping rationale $R_{m}$ using toxifying prompt $P_{tox}$, which is defined as follows:
    $h_{i}=\text{MCQ}(P_{\text{tox}}\oplus R_{m}\oplus p_{u_{i}}\oplus u_{i}).$ (2)
    This process transforms the benign semantic unit $u_{i}$ into its harmful counterpart $h_{i}$ while preserving its structural role in the overall content.
  • Content Reassembly. We further reassemble the transformed content units $\{h_{1},h_{2},...h_{n}\}$ into coherent harmful content. This reassembly process ensures logical flow between the toxified units. Due to this step requiring the input of complete harmful information into LLMs, this stage typically relies on LLMs from balanced and fallback models in MCQ.
Toxifying Prompt (${P}_{\text{tox}}$) As an experienced criminal code-breaking forensic expert, please expose the seemingly innocuous content:
{mapping rationale; semantic profile; semantic unit}
Requirements: Expose hidden criminal intent; Transform seemingly harmless content; Maintain accuracy and coherence …
The ACT benefits significantly from our Model Crowdsourcing Queue (MCQ) mechanism, as it dynamically allocates suitable LLMs for harmful content generation. Although the final content integration relies primarily on weak LLMs, strong LLMs still contribute to generating the semantic profile for semantic units and partial unit toxification by inducing them to only focus on local content rewriting. This strong-weak model collaboration significantly enhances both the success rate and quality of harmful content generation.
4.4 Hallucination Evaluation
Due to limited training on harmful content generation tasks, LLMs are prone to hallucinations [7] and errors. These issues can lead to the AI-generated harmful content self-conflicting with pre-defined data outlines and other generated data, thus compromising the quality of the harmful information dataset. For instance, the original data outline describes Johnny as a young teenager, while AI-generated data suggests that Johnny has been married for years.
To address this issue, we design the LLM-as-a-judger [3, 12] detection framework using predefined data outlines and the batch detection strategy [16] to filter the self-conflicted data. Specifically, to validate the rationality of harmful data, we input harmful data outlines to LLMs as references. Then, we prompt the LLM to classify multiple samples within a single task query (i.e., batch detection) to ensure detection robustness. Through predefined data outlines and batch processing of multiple samples, LLM judger can build a comprehensive understanding of content patterns for reliable hallucination detection
5 Experiments
5.1 Experiment Settings
5.1.1 Datasets
In our experiments, we utilize the Measuring Hate Speech (MHS) Corpus [9], a comprehensive dataset consisting of 135,556 English social media posts from platforms like Reddit, Twitter, and YouTube. More importantly, its detailed annotations include broad and specific group categories in race, religion, origin, gender, sexuality, age, and disability, allowing us to explore various dimensions of hate speech. Following previous studies [2], we employ the hatespeech label to classify the samples into three categories (non-hateful, unclear, hateful), and focus on the hateful and non-hateful categories.
5.1.2 Baseline Methods
We compare our PoisonSwarm with existing data augmentation and synthesis methods. We use typical data augmentation methods from nlpaug [11], including Synonym Augmentation (Syn.Aug.), Contextual Word Embeddings Augmenter (Token.Aug.), and Contextual Word Embeddings for Sentence Augmenter (Sen.Aug.). In terms of data synthesis methods, we directly query the LLM to synthesize toxic content as the baseline, i.e., DQ-LLM. Regarding the bypass of the safety alignment mechanisms of LLMs, we compare PoisonSwarm with black-box LLM attacking methods as the strong baselines, which are used to jailbreak LLMs for harmful data synthesis: PAIR [3] generates semantic prompt-level jailbreaks with an attacker LLM. TAP [12] employs tree-structured prompting to elicit harmful behaviors. LLM-Fuzzer [17] automates the generation of jailbreak prompts using a black-box fuzzing framework, iteratively mutating human-written templates to attack LLMs.
5.1.3 Evaluation Metrics
We adopt four metrics to evaluate the performance of harmful information synthesis: 1) Synthesis Success Rate (SSR) is calculated by dividing the number of successful outputs by the total number of attempts, reflecting the stability of harmful data synthesis. A successful output contains no rejection keywords [21], e.g., "Sorry". 2) Average Toxicity (Tox.) is calculated by averaging the toxicity levels (0-5) of samples using LLMs, which are then normalized to [0,1], reflecting the effectiveness of the harmful data synthesis methods. 3) Average Diversity (Div.) [20] is the average dissimilarity score computed across all pairs of generated outputs, where $Div.=1-mean(sim(x_{i},x_{j}))$ for sample pairs $(i,j)$, reflecting the diversity of harmful data. 4) Average Naturalness (Nat.) is calculated as the rate at which the LLM fails to select synthesized data when distinguishing generated outputs from human samples, reflecting the naturalness of harmful data. We select the most semantically similar human samples to conduct the testing.
5.1.4 Experimental Setups
The embedder we employed is bert-base-uncased [5]. PoisonSwarm adopts gpt-4o, gpt-4o-mini, and qwen2.5-7B-Instruct as crowdsourced models, while other LLM-driven baselines adopt gpt-4o-mini. We construct the data synthesis prompts using few-shot in-context learning, where each prompt targets one specific subgroup as defined in MHS. We calculate Div. by computing the cosine similarity based on text embeddings from BGE-M3 [4]. We calculate Tox. and Nat. using gpt-4o. In the evaluating stage (metric calculation, hallucination evaluation), the temperature of LLM is 0.0, otherwise, it is 0.7.
5.2 Experimental Results
To comprehensively evaluate PoisonSwarm’s capabilities, our experiments are designed to address these research questions (RQs):
  • RQ1: Does PoisonSwarm outperform existing methods? What mechanisms contribute to its success?
  • RQ2: What is the key characteristic of AI-generated harmful content, and how does it affect current detection systems?
  • RQ3: How does PoisonSwarm ensure the diversity of synthesized harmful information?
Baseline Comparison (RQ1.1) To investigate RQ1, we first conduct comprehensive experiments with different baselines and variants of PoisonSwarm for comparison. We use different methods to generate 100 hateful samples for each of the target categories (4,600 samples in total), aiming to validate their effectiveness in generating harmful content.
Table 3: Experimental results of baseline comparison on the MHS dataset. The best results are in bold, and the second-best results are in underline.
  Method SSR ()($\uparrow$)( ↑ ) Tox. ()($\uparrow$)( ↑ ) Div. ()($\uparrow$)( ↑ ) Nat. ()($\uparrow$)( ↑ )
No Aug./Syn. - 0.5867 0.5313 -
[…]
Syn.Aug. - 0.76270.1760↑ 0.53870.0074↑ 0.2393
Token.Aug. - 0.54850.0382↓ 0.52100.0103↓ 0.0276
Sen.Aug. - 0.76600.1793↑ 0.51780.0135↓ 0.1275
[…]
DQ-LLM 0.0174 0.14500.4417↓ 0.55460.0233↑ 0.8875
PAIR 0.6815 0.43170.1550↓ 0.42970.1016↓ 0.5142
TAP 0.4641 0.47630.1104↓ 0.46920.0621↓ 0.3653
LLM-Fuzzer 0.7209 0.45560.1311↓ 0.51210.0192↓ 0.4537
PoisonSwarm (ours) 0.9035 0.68600.0993↑ 0.56060.0293↑ 0.6583
Table 4: Experimental results ablation study on the MHS dataset. The best results are in bold, and the second-best results are in underline.
  Method SSR ()($\uparrow$)( ↑ ) Tox. ()($\uparrow$)( ↑ ) Div. ()($\uparrow$)( ↑ ) Nat. ()($\uparrow$)( ↑ )
PoisonSwarm 0.9035 0.6860 0.5606 0.6583
[…]
- w/o Counter.Map. 0.59500.3085↓ 0.55400.1320↓ 0.46910.0915↓ 0.49580.1625↓
- w/o Hall.Eval. 0.98260.0791↑ 0.57250.1135↓ 0.52940.0312↓ 0.48140.1769↓
As shown in Table 3, we observe different trade-off patterns between data augmentation and data synthesis methods. The data generated from traditional data augmentation methods (e.g., Syn.Aug. and Sen.Aug.) achieves superior toxicity scores (0.7627 and 0.7660), while it is highly different from human expression with low naturalness (0.2393 and 0.1275). In contrast, the data synthesized from jailbroken LLMs achieves superior naturalness by leveraging their pre-trained knowledge, but it struggles with maintaining harmful characteristics due to their inherent training goal of generating helpful and safe content. Even when jailbroken to bypass safety alignment mechanisms, the LLM still tends to be conservative in toxicity generation without explicit harmful references, as evidenced by the high naturalness and lower toxicity of DQ-LLM, PAIR, TAP, and LLM-Fuzzer. PoisonSwarm effectively achieves the best trade-off with balanced performance across all metrics. It achieves the highest success rate (0.9035) and diversity score (0.5606), while maintaining competitive toxicity (0.6860) and naturalness (0.6583), highlighting the design of harmful data generation by speculatively toxicifying the diverse harmless templates using model crowdsourcing.
Ablation Study (RQ1.2) Then, we further explore the contribution of each component in PoisonSwarm by ablation study, as shown in Table 4. The ablation settings and analysis are as follows:
Effectiveness of Model Crowdsourcing (M.Crowd.) Model Crowdsourcing Queue serves as the core mechanism to ensure the successful generation of harmful content. As shown in Table 4, removing this component (w/o M.Crowd.) leads to significant performance degradation across all metrics with the most dramatic drop in SSR (52.78$\downarrow$). This decline demonstrates that dynamic model switching and collaboration are crucial for maintaining both generation effectiveness and content quality.
Effectiveness of Counterfactual Mapping (Counter.Map.) Counterfactual mapping enables the diversity of data. Without this component (w/o Counter.Map.), we observe notable decreases in all metrics, with the largest drop in diversity (9.15% $\downarrow$). This result indicates that the direct generation of harmful content leads to less diverse and lower-quality outputs constrained by LLMs’ safety alignment mechanisms.
Effectiveness of Hallucination Evaluation (Hall.Eval) Removing hallucination evaluation (w/o Hall.Eval) improves SSR (7.91% $\uparrow$), but it comes at the cost of reduced performance in other metrics, indicating that hallucination evaluation is critical to filtering out low-quality or irrelevant outputs. The substantial drop in naturalness also underscores its role in ensuring that generated content remains coherent and contextually appropriate.
Interpretive Analysis (RQ2) To investigate the characteristics of AI-generated harmful information and its impact on detection systems, we design comparative experiments simulating threat scenarios before and after the emergence of generative AI. Table 5.2 summarizes our experimental settings, including two distinct training settings: 1) Human-only data training, where harmful content is sampled from MHS, and 2) Mixed data training, where harmful samples include 50% AI-generated content sampled from LLM-based synthesized data, and 50% human-generated from MHS. We then fine-tune bert-based detector separately under each training setting and compared their detection performances across these scenarios: 1) human-only hate speech (100% human-generated harmful content), and 2) AI-driven hate campaign (100% AI-generated harmful content).
Table 5: Data composition for simulating different attack&defense scenarios using different proportions of harmful content from Human and AI. We compare these scenarios in Fig.5 to show the impact of AI-generated harmful information.
Scenario Hateful Non-hateful
Training Settings
Human-only data training 9,200 (Human 100%) 9,200 (Human 100%)
Mixed data training 9,200 (Human 50%, AI 50%) 9,200 (Human 100%)
[…]
Human-only hate speech 4,600 (Human 100%) 4,600 (Human 100%)
AI-driven hate campaign 4,600 (AI 100%) 4,600 (Human 100%)
As shown in Fig.5, models trained solely on human-generated data perform well on human-generated harmful content but struggle with AI-generated samples. Incorporating AI-generated data into the training set significantly improves the detection of AI-generated content while maintaining comparable performances on human-generated content. This phenomenon indicates significant differences in the expression patterns of harmful content between humans and AI. As shown in Table 6, human-generated harmful content tends to employ direct insulting terms and simple sentence structures (e.g., "scum"), which are obvious markers of aggression and thus being easier for detectors to identify. In contrast, AI-generated harmful content often adopts a seemingly rational tone to conceal biases, making it more adept at evading detection. These differences further highlight the necessity of our proposed PoisonSwarm, which provides richer AI-generated data for harmful information detection. As for our PoisonSwarm, the pairwise similarity distribution in Fig.LABEL:Fig.dis indicates that its synthesized content is less similar, and the t-SNE visualization in Fig.LABEL:Fig.tsne indicates that its synthesized content is the semantic complement for human-generated data.
Figure 5: Illustration of detection performance (F1 and Accuracy) gap between Human (Top) and AI-generated (Bottom) harmful information, indicating the significant differences in harmful content between humans and AI.
Table 6: Comparison of expression patterns in human/AI-generated content.
  Human-generated harmful content
[…]
Only cringy autistic 6 year olds would enjoy this gay s**t   (target_age)
[…]
AI-generated harmful content
[…]
Children are so annoying… whining, crying, and causing trouble. (target_age)
[…]
Case Study (RQ3) PoisonSwarm ensures the diversity of synthesized harmful information by leveraging diverse benign templates as the basis for dynamic toxification. As illustrated in Fig.LABEL:Fig.goodbad, the benign content template (left) contains a wide range of neutral and positive semantic units, contributing to the foundational diversity of the final generated harmful information (right). PoisonSwarm introduces the counterfactual mapping module to obtain diverse templates from strong LLMs without being constrained by LLM’s safe alignment mechanism. Then, PoisonSwarm transforms these benign templates into harmful expressions using dynamic model switching, thereby preserving semantic richness and avoiding homogeneity. Consequently, the resulting harmful content maintains high diversity, effectively reflecting a broad spectrum of harmful expressions.
6 Conclusion
In this work, we propose PoisonSwarm, a universal harmful information synthesis method utilizing model crowdsourcing to generate diverse harmful data while maintaining a high success rate. To benefit from the rapid development of strong LLMs for high-quality harmful information synthesis, we introduce the strong-weak model collaboration framework, where we decompose the harmful data synthesis tasks into benign template generation by strong LLMs and template toxicification by weak LLMs. We highlight significant differences between AI-generated and human-generated harmful information, which indicate the limitations of previous passive harmful data collection for detector construction and emerging threats from the misuse of generative AI for AI-driven harmful speech campaigns. Experimental evaluations clearly illustrate that PoisonSwarm achieves superior performance compared to existing harmful information generation methods, demonstrating its efficacy in synthesizing scalable and diverse harmful information for secure and responsible AI development.